# Robust Control Lyapunov-Value Functions for Nonlinear Disturbed Systems

### v2

@@ Abstract
This article presents a method to construct robust control Lyapunov value functions (R-CLVFs) for robust and stabilizing control of nonlinear systems with input and disturbance bounds. Through modifications to Hamilton-Jacobi reachability analysis, the R-CLVF can be computed via dynamic programming. We prove that the R-CLVF can be used to stabilize the system to its smallest robust control invariant set at a user-specified exponential rate. The R-CLVF additionally can be used to obtain the domain over which stabilization at the desired rate is possible, i.e., the region of exponential stabilizability. Due to the computationally expensive nature of dynamic programming, we additionally propose methods to improve tractability via decomposing the dynamical model or warm-starting the computation. Under certain assumptions, we show that these approaches maintain exact solutions. Three numerical examples are provided, validating our definition of the smallest robust control invariant set, illustrating the impact of the exponential rate and different loss functions, and demonstrating the efficiency of computation using warm-starting and decomposition.
@@ I INTRODUCTION
Standard HJ reachability analysis focuses on problems such as the minimum time to reach a goal, or how to avoid certain states for a finite (or infinite) time horizon. It does not stabilize a system to a goal after reaching it. In our previous work [22], we modified the value function and defined the control Lyapunov value function (CLVF) for undisturbed nonlinear systems with bounded control. The CLVF finds the smallest control invariant set (SCIS) and the region of exponential stabilizability (ROES) of the system given a user-specified exponential rate $\gamma$. Its gradient can be used to synthesize controllers that stabilize the system to the SCIS at the rate $\gamma$. The stabilizability comes from the fact that the CLVF is the viscosity solution to a modified variational inequality. However, there is one technical error in the previous work: the CLVF is not the unique viscosity solution to the VI, and we provide a counterexample in this article. The original work [22] is restricted in several aspects: it considers only systems without disturbance, the choice of the loss function is restricted to 2-norm, the curse of dimensionality restricts its application to relatively low-dimensional systems (5D or lower), and some concepts would benefit from deeper discussion.
Other works that study the relationship between CLF and some partial differential equations (PDE) include Zubov’s method and its extensions [23, 8, 24, 25]. Since the solution to the PDE is generally not continuously differentiable, the differential inclusion is usually considered. Another work that considers the relation between the state-constrained optimal control problem and the HJI-VI is [26], where an auxiliary unconstrained optimal control problem is solved using the HJI-VI. However, this requires augmenting to one additional dimension, which requires much larger computational resources.
This work seeks to address all the restrictions of the previous CLVF work, and provide more rigorous proofs of the theorems. The main contributions are:
@@
    We define the time-varying robust CLVF (TV-R-CLVF) and the robust CLVF (R-CLVF) for systems with bounded disturbance and control. We prove that the R-CLVF is Lipschitz continuous, satisfies the dynamic programming principle (DPP), and is the viscosity solution to the corresponding R-CLVF variational inequality (VI). The algorithm for computing the R-CLVF is updated.
@@
    We define the smallest robustly control invariant set (SRCIS) of a point of interest (POI) given system dynamics and prove that the SRCIS is the zero sub-level set of the R-CLVF. We show that the R-CLVF can stabilize the system to the SRCIS of the POI from the ROES with an exponential rate $\gamma$ defined by the user.
@@
    Two methods to accelerate computation are introduced: warm-starting and system decomposition. We prove that under certain assumptions, these acceleration methods will recover exact results.
@@
    A point-wise optimal R-CLVF quadratic program (QP) control law is provided, which is guaranteed to be feasible, and numerical examples are provided to validate the theory and show numerical efficiency with warm-starting and system decomposition.
The paper is organized in the following order: Section II provides background information on HJ reachability analysis and CLVF. Section III introduces the TV-R-CLVF, and builds up the theoretical foundation for the R-CLVF. A feasibility-guaranteed quadratic program controller is provided. Section IV introduces warm-starting and system decomposition to accelerate the computation. Section V shows three numerical examples, validating the theory.
@@ II BACKGROUND
In this paper, we seek to exponentially stabilize a given nonlinear time-invariant dynamic system with bounded control and disturbance to its SRCIS. We start by defining crucial some terms.
@@ II-A Problem Formulation
where $t<0$ is the initial time, and $x_{0}\in\mathbb{R}^{n}$ is the initial state. The control signal $u(\cdot)$ and disturbance signal $d(\cdot)$ are drawn from the set of measurable functions $\mathbb{U}$ and $\mathbb{D}$. Assume also the control input $u$ and disturbance $d$ are drawn from convex compact sets $\mathcal{U}\subset\mathbb{R}^{m}$ and $\mathcal{D}\subset\mathbb{R}^{p}$ respectively. We have:
$\mathbb{U}:=\{u(\cdot):[t,0]\mapsto\mathcal{U},u(\cdot)\text{ is measurable}\},$
$\mathbb{D}:=\{d(\cdot):[t,0]\mapsto\mathcal{D},d(\cdot)\text{ is measurable}\}.$
@@
Under these assumptions, given an initial state $x$ and control and disturbance signals $u(\cdot)$, $d(\cdot)$, there exists a unique solution $\xi(s;t,x,u(\cdot),d(\cdot))$, $s\in[t,0]$ of the system (1). We call this solution the trajectory of the system, and where possible, we use $\xi(s)$ for conciseness. Further assume the disturbance signal can be determined as a strategy with respect to the control signal: $\lambda:\mathbb{U}\mapsto\mathbb{D}$, drawn from the set of non-anticipative maps $\lambda\in\Lambda$  [27], defined as:
[…]
@@
(RCIS) A closed set $\mathcal{I}$ is robustly control invariant for (1) if $\forall x\in\mathcal{I}$, $\forall\lambda\in\Lambda$, $\exists u(\cdot)\in\mathbb{U}$ such that $\xi(s;t,x,u(\cdot),\lambda[u])\in\mathcal{I}$, $\forall s\in[t,0]$.
@@
    When the system has equilibrium points, the origin 0 is one, i.e. $f($0,0,0)=)=) = 0.
@@
$A_{4}$ is a weaker assumption compared to assuming the existence of a CLF. For a valid CLF, the region of null controllability (RONC) is also robust control invariant. Further, all trajectories starting from the RONC can be stabilized to the equilibrium point. This is considered as one key contribution of the paper: we extend the classical definition of stabilizing to the equilibrium point to stabilizing to the SRCIS, if the former is impossible to achieve.
@@
$\text{dst}(x;\mathcal{A})=\begin{cases}\min_{a\in\partial\mathcal{A}}||x-a||&x\notin\mathcal{A},\\ -\min_{a\in\partial\mathcal{A}}||x-a||&x\in\text{int}(\mathcal{A}),\\ 0&x\in\partial\mathcal{A}.\end{cases}$ (2)
@@
[…]
@@ II-B HJ Reachability and SRCIS
Traditionally in HJ reachability analysis, a continuous loss function $h:\mathbb{R}^{n}\mapsto\mathbb{R}$ is defined such that its zero super-level set is the failure set =$\mathcal{F}=\{x:h(x)>0\}$. The finite-time horizon cost function captures whether a trajectory enters $\mathcal{F}$ at any time in $[t,0]$ under given control and disturbance signals by computing the maximum loss accrued over time:
@@
$V(x,t)$ $=\sup_{\lambda\in\Lambda}\inf_{u(\cdot)\in\mathbb{U}}J(t,x,u(\cdot),\lambda[u])$
$=\sup_{\lambda\in\Lambda}\inf_{u(\cdot)\in\mathbb{U}}\max_{s\in[t,0]}h(\xi(s;t,x,u(\cdot),\lambda[u]).$ (4)
@@
Different from the time-varying value function (II-B), for all states in the $\alpha$ sub-level set of $V^{\infty}(x)$, there always exists a control signal such that the maximum loss is lower than $\alpha$ despite the disturbance signal. This means every $\alpha$ sub-level set of $V^{\infty}(x)$ is robustly control invariant, and the trajectories can be maintained within a particular level set boundary. Further, this set is the largest RCIS contained within the $\alpha$ sub-level set of $h(x)$ [29].
@@
In this paper, we restrict the selection of $h(x)$ to be vector norms (e.g., p-norms, or weighted Q norms,) and the same as the norm used in equation (2). More specifically, given any POI $p\in\mathbb{R}^{n}$, we pick $h(x;p)=\|x-p\|$. In other words, the loss function measures the distance of a state to the POI. With this restriction, the cost function (3) captures the largest deviation from the POI of a given trajectory, initialized at $x$ with $u(\cdot)$ and $d(\cdot)$ applied, in time horizon $[t,0]$. The infinite time value function (5) captures the largest deviation with optimal control and disturbance signals applied in an infinite time horizon.
Denote the minimal value of $V^{\infty}(x)$ as $V^{\infty}_{m}:=\min_{x}V^{\infty}(x)$. The $V^{\infty}_{m}$-level set of $V^{\infty}$ is the smallest RCIS (SRCIS), and denoted by $\mathcal{I}_{m}$. Different norms result in different SRCISs.
@@
(SRCIS.) A closed set $\mathcal{I}_{m}(p)$ is the SRCIS of a given POI $p$ and norm chosen if it is the level set corresponding to the minimal value of $V^{\infty}$.
Since the SRCIS is the level set corresponding to the minimal value, all states in it have the same value $V^{\infty}_{m}$, and we have
@@
where $u\in[-1,1]$ and $d\in[-0.5,0.5]$. This system has an undisturbed, uncontrolled equilibrium point $[x,u,d]=[\textbf{0},0,0]$. It can be verified that =$\mathcal{I}=\{x_{1}\in[-0.5,0.5],x_{2}=0\}$ is one ‘minimal RCIS’ as all its subsets are not robustly control invariant. In fact, picking any $x_{2}\in[-1,1]$ results in a ‘minimal RCIS.’ On the other hand, picking $h(x)=||x||_{\infty}$, the SRCIS is $\mathcal{I}_{m}=\{x_{1},x_{2}\in[-0.5,0.5]\}$. This is because though the control can stabilize any $|x_{2}|<1$ to the origin, the disturbance is also strong enough to perturb any $|x_{1}|<0.5$ to leave the origin. Therefore, all states s.t. $x_{1},x_{2}\in[-0.5,0.5]$ have the same value, and the SRCIS measured by the $\infty$-norm is a square. Fig. 1 shows the SRCIS for three different choices of $h(x)$ and the corresponding value function.
Figure 1: SRCIS corresponds to different loss functions for system (8). Top left to right: different loss functions, including the 2-norm, infinity norm, and weighted Q-norm ($Q=diag[0.2,1]$). Middle left to right: R-CLVF ($\gamma=0$) when $h(x)=||x||_{2}$, $||x||_{\infty}$, $||x||_{Q}$, and $||x||_{Q}=\sqrt{x^{T}Qx}$. Bottom left to right: the corresponding SRCIS and a trajectory starting inside the SRCIS. The robust control invariance is validated.
@@
Proposition 1
Define $\underline{h}(x)=h(x)-a$, and denote the corresponding value function as $\underline{V}(x,t)$, we have $\underline{V}(x,t)=V(x,t)-a$.
Proof:
$\underline{V}(x,t)$ $=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max_{s\in[t,0]}\underline{h}(\xi(s;t,x,u(\cdot),\lambda[u])$
$=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max_{s\in[t,0]}\big{(}h(\xi(s;t,x,u(\cdot),\lambda[u])-a\big{)}$
$=V(x,t)-a,$ (9)
where the last step is because $a$ is a constant. ∎
Also note that we dropped $p$ from $h$ in Prop. 1, because it does not affect the result.
@@ III ROBUST CONTROL LYAPUNOV-VALUE FUNCTIONS
In this section, we start by defining the time-varying robust CLVF (TV-R-CLVF) and prove some important properties of it. We then define the R-CLVF, which is the limit function of the TV-R-CLVF. We show that the existence of the R-CLVF is equivalent to the exponential stabilizability of the system to its SRCIS of a POI and that its domain is the ROES. This section is organized as follows: Section III-A defines the TV-R-CLVF and shows some key properties of it. Section III-B defines the R-CLVF and shows it satisfies the DPP (Theorem 3) and is the viscosity solution of a variational inequality (Theorem 4). Section III-C shows the sufficient and necessary conditions for the existence of the R-CLVF (Theorem 7). Section III-D studies the impact of different $\gamma$. It shows for all non-negative $\gamma$, the SRCISs are all the same (Lemma 8). Further, a larger $\gamma$ corresponds to a faster convergence rate, while a smaller ROES. Section III-E provides an feasibility guaranteed QP controller.

### v3

@@ Abstract
This article presents a method to construct robust control Lyapunov value functions (R-CLVFs) for robust and stabilizing control of nonlinear systems with input and disturbance bounds. Through modifications to Hamilton-Jacobi reachability analysis, the R-CLVF can be computed via dynamic programming. We prove that the R-CLVF can be used to stabilize the system to the smallest robust control invariant set of any point of interest at a user-specified exponential rate. The R-CLVF additionally can be used to obtain the domain over which stabilization at the desired rate is possible, i.e., the region of exponential stabilizability. Due to the computationally expensive nature of dynamic programming, we additionally propose methods to improve tractability via decomposing the dynamical model or warm-starting the computation. Under certain assumptions, we show that these approaches maintain exact solutions. Three numerical examples are provided, validating our definition of the smallest robust control invariant set, illustrating the impact of the exponential rate and different loss functions, and demonstrating the efficiency of computation using warm-starting and decomposition.
Keywords: Nonlinear systems, Optimal control, Robust control, Stability of nonlinear systems.
@@ I INTRODUCTION
Standard HJ reachability analysis focuses on problems such as the minimum time to reach a goal, or how to avoid certain states for a finite (or infinite) time horizon. It does not stabilize a system to a goal after reaching it. In our previous work [29], we modified the value function and defined the control Lyapunov value function (CLVF) for undisturbed nonlinear systems with compact control. The CLVF finds the smallest control invariant set and the region of exponential stabilizability (ROES) of the system given a user-specified exponential rate $\gamma$. The original work is restrictive and suffers from the curse of dimensionality.
Other works that study the relationship between CLF and some partial differential equations (PDE) include Zubov’s method and its extensions [43, 30, 20, 10][43] is the first that relates the Lyapunov function with a PDE, called Zubov’s equation. Later, [30, 20, 10] extends this early work to systems with control. The general process is to define an optimal control problem, whose value function itself is a CLF and is the viscosity solution to some PDE of Zubov’s type. Therefore, the CLF (which is the value function) can be found by solving the PDE. Viscosity solutions are considered because these PDEs usually do not admit smooth solutions. Another work that considers the relation between the state-constrained optimal control problem and the HJI-VI is [1], where an auxiliary unconstrained optimal control problem is solved using the HJI-VI. However, this requires augmenting to one additional dimension, which requires much larger computational resources. Further, there also exist works on computing the region of attraction of a given system. These methods use optimization like linear matrix inequality [14], sum of squares [40, 39], or convex program [33]. However, they usually only guarantee under approximation and require special forms of the dynamics (e.g., polynomial dynamics). In contrast, we provide exact recovery of the ROES for general nonlinear systems, and we consider the convergence to a set, instead of just the equilibrium point.
This work seeks to extend our previous CLVF work, correct the errors, and provided numerical feasible approach for high-dimensional systems. The main contributions are:
@@
    We define the smallest robustly control invariant set (SRCIS) of a point of interest (POI) given system dynamics. We extend our previous work [29] and define the robust-CLVF (R-CLVF), which is with respect to an arbitrary POI, instead of an equilibrium point.
@@
    We prove that the R-CLVF is Lipschitz continuous, satisfies the DP principle (DPP), is a viscosity solution to the corresponding R-CLVF variational inequality (VI), and we update the algorithm for computing it. We also show that the SRCIS is the zero sub-level set of the R-CLVF, and that the R-CLVF can stabilize the system to the SRCIS of the POI from the ROES with an exponential rate $\gamma$ defined by the user. We also show that the domain of the R-CLVF is exactly the ROES.
@@
    Our method uses $\gamma$ as an exponential amplifier, not as a discount factor (which is common in the literature). This amplifier provides straightforward physical intuition for our formulation.
@@
    Two methods to accelerate computation are introduced: warm-starting and system decomposition. We prove that under certain assumptions, these acceleration methods will recover exact R-CLVF.
The paper is organized in the following order: Section II provides background information on HJ reachability analysis. Section III introduces the R-CLVF and presents the main results of the paper, e.g., the existence of the R-CLVF is equivalent to the exponential stabilizability of the system. A feasibility-guaranteed quadratic program (QP) controller is provided. Section IV introduces warm-starting and system decomposition to accelerate the computation. Section V shows three numerical examples, validating the theory.
@@ II BACKGROUND
In this paper, we seek to exponentially stabilize a given nonlinear time-invariant dynamic system with compact control and disturbance to its SRCIS. We start with the background information.
@@ II-A Problem Formulation
where $t<0$ is the initial time, and $x_{0}\in\mathbb{R}^{n}$ is the initial state. The control signal $u(\cdot)$ and disturbance signal $d(\cdot)$ are drawn from the set of measurable functions $\mathbb{U}_{t}$ and $\mathbb{D}_{t}$. Assume also the control input $u$ and disturbance $d$ are drawn from convex compact sets $\mathcal{U}\subset\mathbb{R}^{m}$ and $\mathcal{D}\subset\mathbb{R}^{p}$ respectively. We have:
$\mathbb{U}_{t}:=\{u(\cdot):[t,0]\mapsto\mathcal{U},u(\cdot)\text{ is measurable}\},$
$\mathbb{D}_{t}:=\{d(\cdot):[t,0]\mapsto\mathcal{D},d(\cdot)\text{ is measurable}\}.$
@@
Under these assumptions, given an initial state $x$ and control and disturbance signals $u(\cdot)$, $d(\cdot)$, there exists a unique solution $\xi(s;t,x,u(\cdot),d(\cdot))$, $s\in[t,0]$ of the system (1). We call this solution the trajectory of the system, and where possible, we use $\xi(s)$ for conciseness. Further assume the disturbance signal can be determined as a strategy with respect to the control signal: $\lambda:\mathbb{U}_{t}\mapsto\mathbb{D}_{t}$, drawn from the set of non-anticipative maps $\lambda\in\Lambda_{t}$  [41], defined as:
$\Lambda_{t}:=\{$ $\mathcal{N}:A(t)\mapsto B(t):a(s)=\hat{a}(s)\text{ a.e. }\forall s\in[t,0]$
@@
(RCIS) A closed set $\mathcal{I}$ is robustly control invariant for (1) if $\forall x\in\mathcal{I}$, $\forall\lambda\in\Lambda_{t}$, $\exists u(\cdot)\in\mathbb{U}_{t}$ such that $\xi(s;t,x,u(\cdot),\lambda[u])\in\mathcal{I}$, $\forall s\in[t,0]$.
@@
    When the system has equilibrium points, the origin 0 is one, i.e. f(f(0,0,0)=)= 0.
@@
$A_{4}$ is a weaker assumption compared to assuming the existence of a CLF. For a valid CLF, the region of null controllability (RONC) is also robust control invariant. Further, all trajectories starting from the RONC can be stabilized to the equilibrium point. With $A_{4}$, we extend the classical definition of stabilizing to the equilibrium point to stabilizing to the SRCIS, if the former is impossible to achieve.
@@
[…]
@@
$\mathcal{D}_{\text{ROES}}:=\{x\in\mathbb{R}^{n}|\hskip 5.69054pt\forall\lambda\in\Lambda_{t},\exists u(\cdot)\in\mathbb{U}_{t},\gamma,c>0\text{ s.t. }$
@@ II-B HJ Reachability and SRCIS
Traditionally in HJ reachability analysis, a continuous loss function $h:\mathbb{R}^{n}\mapsto\mathbb{R}$ is defined such that its zero super-level set is the failure set $\mathcal{F}=\{x:h(x)>0\}$. The finite-time horizon cost function captures whether a trajectory enters $\mathcal{F}$ at any time in $[t,0]$ under given control and disturbance signals by computing the maximum loss accrued over time:
@@
$V(x,t)$ $=\sup_{\lambda\in\Lambda_{t}}\inf_{u(\cdot)\in\mathbb{U}_{t}}J(t,x,u(\cdot),\lambda[u]),$
$=\sup_{\lambda\in\Lambda_{t}}\inf_{u(\cdot)\in\mathbb{U}_{t}}\max_{s\in[t,0]}h(\xi(s;t,x,u(\cdot),\lambda[u]).$ (4)
@@
Different from the time-varying value function (II-B), for all states in the $\alpha$ sub-level set of $V^{\infty}(x)$, there always exists a control signal such that the maximum loss is lower than or equal to $\alpha$ despite the disturbance signal. This means every $\alpha$ sub-level set of $V^{\infty}(x)$ is robustly control invariant, and the trajectories can be maintained within a particular level set boundary. Further, this set is the largest RCIS contained within the $\alpha$ sub-level set of $h(x)$ [15].
@@
In this paper, we restrict the selection of $h(x)$ to be vector norms (e.g., p-norms, or weighted Q norms), and specifically, it is the same norm used in equation (2). More specifically, given any POI $p\in\mathbb{R}^{n}$, we pick $h(x;p)=\|x-p\|$ (where $p$ is a hyperparameter). In other words, the loss function measures the distance of a state to the POI. With this restriction, the cost function (3) captures the largest deviation from the POI of a given trajectory, initialized at $x$ with $u(\cdot)$ and $d(\cdot)$ applied, in time horizon $[t,0]$. The infinite time value function (5) captures the largest deviation with optimal control and disturbance signals applied in an infinite time horizon. Further, the limit in (5) may not exist globally, but always exists on any compact subset of $\mathbb{R}^{n}$ under our assumptions.
@@
(SRCIS.) Given a POI $p$, a norm, compute the value function $V^{\infty}(x)$ in (5) and denote its minimal value as $V^{\infty}_{m}=\min_{x}V^{\infty}(x)$. The SRCIS, denoted as $\mathcal{I}_{m}$, is the $V^{\infty}_{m}$-level set of $V^{\infty}$.
Note that any $\alpha$-sublevel set of $V^{\infty}$ is closed and bounded, therefore the minimal value $V^{\infty}_{m}$ is always attained. Further, since SRCIS is the level set corresponding to the minimal value, all states in it have the same value $V^{\infty}_{m}$. This also holds for the state $\bar{x}=\operatorname*{argmax}_{x\in\mathcal{I}_{m}}h(x;p)$. Because $\mathcal{I}_{m}$ is robustly control invariant, the trajectory starting from $\bar{x}$ stays in $\mathcal{I}_{m}$, and it’s maximum deviation from $p$ is $\max_{x\in\mathcal{I}_{m}}h(x;p)$. Combined, we have:
@@
[…]
@@
Remark 4
Define $\underline{h}(x)=h(x)-a$, and denote the corresponding value function as $\underline{V}(x,t)$, we have $\underline{V}(x,t)=V(x,t)-a$. This is because
$\underline{V}(x,t)$ $=\sup_{\lambda\in\Lambda_{t}}\inf_{u\in\mathbb{U}_{t}}\max_{s\in[t,0]}\underline{h}(\xi(s;t,x,u(\cdot),\lambda[u]),$
$=\sup_{\lambda\in\Lambda_{t}}\inf_{u\in\mathbb{U}_{t}}\max_{s\in[t,0]}\big(h(\xi(s;t,x,u(\cdot),\lambda[u])-a\big),$
$=V(x,t)-a.$
The last step is because $a$ is a constant. This holds for all $t\leq 0$, so we also have
$\underline{V}^{\infty}(x)=\lim_{t\rightarrow-\infty}\underline{V}(x,t)=V^{\infty}(x)-a.$
@@ III ROBUST CONTROL LYAPUNOV-VALUE FUNCTIONS
In this section, we will pose an optimal control problem, whose value function is a viscosity solution to an HJ-type VI. This VI can be used to prove that the value function has CLF properties, i.e., it implies exponential stabilizability. We highlight that we use an exponential amplifier in our value function instead of a discount factor (which is common in the literature), and that our formulation has very straightforward intuition. This section is organized as follows: Section III-A defines the TV-R-CLVF and shows some key properties of it. Section III-B defines the R-CLVF, shows the sufficient condition for its existence, and shows it satisfies the DPP (Theorem 5) and is the viscosity solution of a VI (Theorem 6). Section III-C shows that the existence of the R-CLVF (Theorem 8) is equivalent to the exponential stabilizability to the SRCIS. Section III-D provides an feasibility guaranteed QP controller.